OpenStudy (anonymous):

find the value of y for the given values of x and z. y=x+z x=6 z=18

Directrix (directrix):

Given equation: y=x+z x = 6 and z = 18

Directrix (directrix):

Substitute 6 for x and 18 for z

Directrix (directrix):

y=x+z y = 6 + 18 y = ?

OpenStudy (anonymous):

24

Directrix (directrix):

Yes.

OpenStudy (anonymous):

wow that was easy!!! I am over-thinking it.